# Homological data on the periodic structure of self-maps on wedge sums

### v1

@@ Abstract.
In this article, we study the periodic points for continuous self-maps on the wedge sum of topological manifolds, exhibiting a particular combinatorial structure. We compute explicitly the Lefschetz numbers, the Dold coefficients and consider its set of algebraic periods. Moreover, we study the special case of maps on the wedge sum of tori, and show some of the homological obstructions present in defining these maps.
@@ 1. Introduction
We are interested in finding certain homological information bounding the number of periodic points and their periods of a continuous self-map on the wedge sum of topological manifolds. We recall that the wedge sums of those spaces are not topological manifolds ([19, p. 10]). One of the simple maps on this category are the graph maps, i.e. self-maps on the wedge sum of circles, which has been extensively studied, in particular its periodic structure, see for instance [1, 16, 17, 22, 27, 28, 31]. Also, self-maps on the wedge sum of spheres has been studied in different contexts [3, 9, 29, 32, 33]. Besides these cases, the dynamics of maps on wedge sums of other spaces has not been studied in detail, as far as we know.
In this article we deal with the calculation of the Lefschetz numbers, the Lefschetz zeta functions, Dold coefficients (the Dold coefficient of order $m$ is also known in the literature as the Lefschetz number of period $m$) and the algebraic periods for general self-maps on wedge sum of compact topological manifolds. We give explicit and closed formulae in term of some more basic maps called the coordinate maps. In particular, Theorem 1 reduces the Lefschetz numbers and the Lefschetz zeta function of the map $f$ in terms of the Lefschetz numbers and zeta functions of its coordinate maps. Theorem 2 deals with the Dold coefficients and Corollary 3 with the algebraic periods for such map.
The study of maps without periodic points has been done in different contexts and spaces, see for instance [4, 18, 20, 23]. A weaker notion is the concept of Lefschetz periodic point free map, i.e. a map whose all Lefschetz numbers are equal to zero, for example see these articles in the subject [13, 15, 29, 33]. In Section 4 we deal about this subject and we prove in Theorem 4, that a continuous permutative self map on the wedge sum of tori is not Lefschetz periodic point free unless the induced maps on homology has zero eigenvalues.
@@
The article is structured as follows in Section 2 we introduced the basic notions and present the main results and statements of the article. In Section 3, we prove the main results and introduce the constructions required in their proofs. In Section 4 we deal with the particular case of self-maps on wedge sum of tori. In Section 5 we give a list of examples that illustrate the theory developed throughout the article and some of the challenges present in the generalization of the context expounded here, which are summarized in the next section. In this Section 6 we provide a list of open problems and conjectures that have arisen in the course of our research, and we also discuss some general comments.
@@ 2. Preliminaries and statements of the main results
If $X={\mathbb{S}}^{1}$ then ${\mathbb{S}}^{1}\vee\cdots\vee{\mathbb{S}}^{1}$ is a graph, in the particular case; ${\mathbb{S}}^{1}\vee{\mathbb{S}}^{1}$ the space is homeomorphic to the figure “8”. In the case of $X={\mathbb{S}}^{2}$ the space ${\mathbb{S}}^{2}\vee\cdots\vee{\mathbb{S}}^{2}$ is called a bouquet of spheres.
Let the maps: $\iota_{j}:X_{j}\to\bigvee_{i=1}^{s}X_{i}$ be the natural inclusion and $\nu_{j}:\bigvee_{i=1}^{s}X_{i}\to X_{j}$ be projection, where $1\leq j\leq s$. Given a continuous map $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$, it induces continuous maps: $f_{j}:\bigvee_{i=1}^{s}X_{i}\to X_{j}$ by $f_{j}=\nu_{j}\circ f$ and $f_{jk}:X_{j}\to X_{k}$ so that: $f_{jk}:=\nu_{k}\circ f\circ\iota_{j}$ and $f_{j}=f_{j1}\vee\cdots\vee f_{js}$ ( for the definition of the wed sum of maps see [34, p. 39]).. The maps $f_{jk}$ are called the the coordinates (or coordinate-maps) of $f$.
Note that $f(x_{0})=x_{0}$ since $f$ is continuous pointed map, where $x_{0}$ is the base point of $\bigvee_{i=1}^{s}X_{i}$.
@@
Let $f$ be a permutative self-map on $\bigvee_{j=1}^{s}X_{j}$ with permutation $\sigma$ and $\sigma=\pi_{1}\cdots\pi_{q}$ be its decomposition into cycles. Let $\Lambda_{1},\ldots,\Lambda_{q}$ be a partition of the index set $\{1,\ldots,s\}$, such that the cardinality of $\Lambda_{j}$ is $s_{j}$, and $s_{1}+\cdots+s_{q}=s$, so that the cycle $\pi_{j}$ acts on the $\Lambda_{j}$ and $f|_{\Lambda_{j}}:\bigvee_{\ell\in\Lambda_{j}}X_{\ell}\to\bigvee_{\ell\in\Lambda_{j}}X_{\ell}$, is a permutative map with permutation $\pi_{j}$. Hence $f=f_{\Lambda_{1}}\vee\cdots\vee f_{\Lambda_{q}}$ is the decomposition described in detailed in Lemma 5 stated in the following section.
@@
We shall require an additional condition on the spaces $X_{i}$ for this class of maps. We say that a permutative self-map on $\bigvee_{i=1}^{s}X_{i}$ is squared by blocks if the spaces $X_{j}$, for $j\in\Lambda_{i}$, $1\leq i\leq q$, are pair wise homeomorphic. In the particular case that the permutative square by blocks map, associated with a cyclic permutation, all the spaces $X_{j}$, for $1\leq j\leq s$ are homeomorphic.
In the following lines we introduce the Lefschetz numbers and the Lefschetz zeta function. Let $Y$ be an Euclidean neighborhood retract (ENR for short) of dimension $N$ (cf. [19, p. 527]). We recall that compact polyhedra, CW complexes are ENR-s, moreover the class of ENR is closed under wedge sums. Let $f:Y\to Y$ be a continuous map, and $f_{*k}:H_{k}(Y,\mathbb{Q})\to H_{k}(Y,\mathbb{Q})$ the induced maps on the $k$-rational homology group of $Y$, with $0\leq k\leq N$. The Lefschetz number $L(f)$ is defined as
@@
A classic result connecting the algebraic topology with the fixed point theory is the Lefschetz Fixed Point Theorem which establishes the existence of a fixed point if $L(f)\neq 0$ (cf. [24, 8]). As it is well-known the converse of this theorem is not true, consider for example the identity map on the torus.
Theorem 1 reduces the Lefschetz numbers and the Lefschetz zeta function of the map $f$ in terms of the Lefschetz numbers and zeta functions of its coordinate maps.
@@
Let $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ be a continuous permutative squared by blocks self-map with associated permutation $\sigma$. Let $s=s_{1}+s_{2}+\cdots+s_{q}$ be the partition associated with the cyclic representation of $\sigma$. Then, for any partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of the set $\{1,2,\ldots,s\}$ for which #$\#\Lambda_{j}=s_{j}$ for each $j=1,2,\ldots,q$, we have that
@@
    $\mathbbm{1}_{s_{l}}(m)=\left\{\begin{array}[]{ll}1&\text{ if }m\equiv 0\pmod{s_{l}}\\ 0&\text{ if }m\not\equiv 0\pmod{s_{l}}.\end{array}\right.$
@@
$\mu(m):=\left\{\begin{array}[]{ll}1&\mbox{ if }m=1,\\ 0&\mbox{ if }k^{2}|m\mbox{ for some }k\in\mathbb{N},\\ (-1)^{r}&\mbox{ if }m=p_{1}\cdots p_{r}\mbox{ has distinct primes factors.}\end{array}\right.$
@@
We remark the fact $\ell(f^{m})\neq 0$ does not necessarily implies the existence of a periodic point of period $m$ for $f$. This is true when the map $f:Y\to Y$ and the space $Y$ has a additional structure. For example when the space $Y$ is a differentiable manifold and $f$ is a Morse-Smale diffeomorphism (see [30], for a survey on the subject) or a $C^{1}$ transversal map (cf. [25]). The terminology of the set of algebraic periods is mentioned explicitly in  [14]. The intersection of the set of algebraic periods and the odd non-negative integers have been widely studied in the context of quasi–unipotent maps (in this case this intersection set is called the minimal set of Lefschetz periods) during the last years, see for example [30] and within references.
@@
Let $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ be a continuous permutative squared by blocks self-map with associated permutation $\sigma$. Let $s=s_{1}+s_{2}+\cdots+s_{q}$ be the partition associated with the cyclic representation of $\sigma$. Then, for any partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of the set $\{1,2,\ldots,s\}$ for which #$\#\Lambda_{j}=s_{j}$ for each $j=1,2,\ldots,q$, we have for $m>1$,
@@
A straight-forward consequence of this theorem is the following corollary:
@@
An interesting application of Theorem 1 in the case of self-maps on the wedge sum of tori is Theorem 4. We recall that a continuous map $f:X\to X$ is Lefschetz periodic point free if $L(f^{m})=0$, for all $m\geq 1$, i.e. $\zeta_{f}(t)\equiv 1$.
@@
Let $f:\bigvee_{i=1}^{s}\mathbb{T}^{n}\to\bigvee_{i=1}^{s}\mathbb{T}^{n}$ be as a continuous permutative map with non-zero eigenvalues them $f$ is not Lefschetz periodic point free.
@@ 3. Proofs of the main results and applications
Throughout the article we shall asume that the spaces $X_{j}$ are path-connected. Given a continuous map $f\colon\bigvee_{j=1}^{s}X_{j}\to\bigvee_{j=1}^{s}X_{j}$ we shall see in the next lemma that it is possible to find a decomposition of the sum $\bigvee_{j=1}^{s}X_{j}$ into subsums of the form $\bigvee_{j\in\Lambda}X_{j}$ which are invariant under the action of $f$, each summand being indecomposable in the sense that no nontrivial subsum has this invariance property. We recall that a set $Z\subset X$ is an invariant set for the map $g\colon X\to X$, if $g(Z)\subset Z$.
@@
Let us consider the graph $\mathcal{G}_{f}=(\mathcal{V}_{f},\mathcal{E}_{f})$ having as a set of vertices $\mathcal{V}_{f}=\{1,2,\ldots,s\}$ and edges defined by the condition $\{i,j\}\in\mathcal{E}_{f}$ if and only if there exist $x_{i}\in X_{i}$ and $x_{j}\in X_{j}$ with $x_{i},x_{j}\not=x_{0}$ (here $x_{0}$ is the base point) such that $f(x_{i})=x_{j}$ or $f(x_{j})=x_{i}$. Next, an equivalent relation is defined by $i\sim_{f}j$ if and only if $i=j$ or there is a path in $\mathcal{G}_{f}$ joining $i$ and $j$. The equivalent classes form the required partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of $\{1,2,\ldots,s\}$. Indeed: the fact $i\in\Lambda_{k}$ and $f(X_{i})\cap X_{j}\not=\emptyset$ implies $j\in\Lambda_{k}$, yields $f(\bigvee_{\ell\in\Lambda_{k}}X_{\ell})\subset\bigvee_{\ell\in\Lambda_{k}}X_{\ell}$. Similarly $f(\bigvee_{\ell\not\in\Lambda_{k}}X_{\ell})\subset\bigvee_{\ell\not\in\Lambda_{k}}X_{\ell}$, the indecomposibility is a consequence of the connectness of the classes of the graph $\mathcal{G}_{f}$. ∎
@@
Lets consider the example of the permutation map described before with $s=3$:
@@
where $O_{{ij}_{*k}}:H_{k}(X_{i},\mathbb{Q})\to H_{k}(X_{j},\mathbb{Q})$ is the null linear map, and $\sigma(1)=2$, $\sigma(2)=1$, $\sigma(3)=3$. In the case that $f$ is squared by blocks, we have $X_{1}\simeq X_{2}$.
@@
In Theorem 1 we express the Lefschetz numbers and the Lefschetz zeta function of a squared by blocks permutative map on $\bigvee_{j=1}^{s}X_{j}$ in terms of its coordinate maps. The strategy of the proof is first to compute them for diagonal maps (Lemma 6), later for cyclic permutative maps (Lemma 7) and finally these results are used for proving the general case.
@@
We would like to point out that the statement (a) of Lemma 6 generalizes Theorem 3.2.1 of [11], in the context of wedge sum of spaces.
@@
$L(f^{m}_{\Lambda_{l}})=\left\{\begin{array}[]{ll}1+\sum_{i\in\Lambda_{l}}\left(L(f^{m}_{ii})-1\right),&m\not\equiv 0\pmod{s_{l}};\\ &\\ 1,&m\not\equiv 0\pmod{s_{l}}.\end{array}\right.$
@@
For the purpose of providing a formula that can be used for effective computations we can consider particular examples of the sets $\Lambda$’s. For example, if the orbits of $\sigma$ are segments of non-decreasing size, then they must begin at the points $1$, $1+s_{1}$, $1+s_{1}+s_{2}$, $\ldots,1+s_{1}+\cdots+s_{q-1}$ and consequently have the form
@@
The Lefschetz zeta function of a continuous self-map $f:Y\to Y$, on any compact polyhedra $Y$, is a rational function. Moreover it satisfies the formula (cf. [12, pg. 38]):
@@
In Theorem 2, we express the Dold coefficients of a squared by blocks permutative map on $\bigvee_{j=1}^{s}X_{j}$ in terms of its coordinate maps. The strategy of the proof is similar to the one used in the proof of Theorem 1, i.e. we consider first the case of diagonal maps (Lemma 10), later we consider cyclic permutative maps (Lemma 11) and finally these results are used for proving the general case.
@@
Part (a) is straightforward from Lemma 7-(a). For part (b), notice that if $r|m$ and $m\not\equiv 0\pmod{s}$ then $r\not\equiv 0\pmod{s}$. Lemma 7-(b), states $L(f^{r})=1$, for $r\not\equiv 0\pmod{s}$. Hence, for $m\not\equiv 0\pmod{s}$, it follows
@@
$\ell(f_{\Lambda_{j}}^{m})=\left\{\begin{array}[]{ll}\displaystyle\sum_{i\in\Lambda_{j}}\ell(f_{ii}^{m}),&\mbox{ if }m\equiv 0\pmod{s_{j}};\\ &\\ 0,&\mbox{ if }m\not\equiv 0\pmod{s_{j}}.\end{array}\right.$
@@
4. Self-maps on wedge sum of tori
In this section we shall consider continuous map on the wedge sum of $n$-dimensional tori, i.e. $\bigvee_{i=1}^{s}\mathbb{T}^{n}$, with $n>1$. As we pointed out in the Introduction, when $n=1$, we have $\bigvee_{i=1}^{s}{\mathbb{S}}^{1}$ is a graph and graph maps have been extensively studied.

### v2

@@ Abstract.
In this article, we study the periodic points for continuous self-maps on a wedge sum of topological manifolds, exhibiting a particular combinatorial structure. We compute explicitly the Lefschetz numbers, the Dold coefficients and consider its set of algebraic periods. Moreover, we study the special case of maps on a wedge sum of tori, and show some of the homological obstructions present in defining these maps.
@@ 1. Introduction
We are interested in finding certain homological information bounding the number of periodic points and their periods for a continuous self-map on a wedge sum of topological manifolds. We recall that the wedge sums of those spaces are not topological manifolds ([19, p. 10]). One class of simple maps on this category are graph maps, i.e. self-maps on a wedge sum of circles, whose periodic point set structure has been extensively studied, see for instance [1, 16, 17, 22, 27, 28, 31]. Also, self-maps on a wedge sum of spheres has been studied in different contexts [3, 9, 29, 32, 33]. Besides these cases, the dynamics of maps on wedge sums of other spaces have not been studied in detail as far as we know.
In this article we deal with the calculation of the Lefschetz numbers, the Lefschetz zeta functions, Dold coefficients (the Dold coefficient of order $m$ is also known in the literature as the Lefschetz number of period $m$) and the algebraic periods for general self-maps on wedge sums of compact topological manifolds. We give explicit and closed formulae in terms of some more basic maps called coordinate maps. In particular, Theorem 1 reduces the Lefschetz numbers and the Lefschetz zeta function of the map $f$ in terms of the Lefschetz numbers and zeta functions of its coordinate maps. Theorem 2 deals with the Dold coefficients and Corollary 3 with the algebraic periods for such map.
The study of maps without periodic points has been done in different contexts and spaces, see for instance [4, 18, 20, 23]. A weaker notion is the concept of Lefschetz periodic point free map, i.e. a map whose all Lefschetz numbers are equal to zero, for example see these articles in the subject [13, 15, 29, 33]. In Section 4 we deal about this subject and we prove in Theorem 4, that establishes that a continuous permutative self map on a wedge sum of tori is not Lefschetz periodic point free unless the induced maps on homology have zero eigenvalues.
@@
The article is structured as follows in Section 2 we introduce the basic notions and present the main results and statements of the article. In Section 3, we prove the main results and introduce the constructions required in their proofs. In Section 4 we deal with the particular case of self-maps on wedge sums of tori. In Section 5 we give a list of examples that illustrate the theory developed throughout the article and some of the challenges present in the generalization of the context expounded here, which are summarized in the next section. In Section 6 we provide a list of open problems and conjectures that have arisen in the course of our research, and we make some general comments.
@@ 2. Preliminaries and statements of the main results
A wedge sum of circles is a graph. In the particular case of two circles, this space is the well-known figure “8". A wedge sum of two-dimensional spheres is called a bouquet of spheres.
Let the maps: $\iota_{j}:X_{j}\to\bigvee_{i=1}^{s}X_{i}$ be the natural inclusion and $\nu_{j}:\bigvee_{i=1}^{s}X_{i}\to X_{j}$ be projection, where $1\leq j\leq s$. Given a continuous map $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$, it induces continuous maps: $f_{j}:\bigvee_{i=1}^{s}X_{i}\to X_{j}$ by $f_{j}=\nu_{j}\circ f$ and $f_{jk}:X_{j}\to X_{k}$ so that: $f_{jk}:=\nu_{k}\circ f\circ\iota_{j}$ and $f_{j}=f_{j1}\vee\cdots\vee f_{js}$ (for the definition of the wedge sum of maps see [34, p. 39]). The maps $f_{jk}$ are called the coordinates (or coordinate-maps) of $f$.
Note that $f(x_{0})=x_{0}$ since $f$ is a continuous pointed map, where $x_{0}$ is the base point of $\bigvee_{i=1}^{s}X_{i}$.
@@
Let $f$ be a permutative self-map on $\bigvee_{j=1}^{s}X_{j}$ with permutation $\sigma$ and $\sigma=\pi_{1}\cdots\pi_{q}$ be its decomposition into cycles. Let $\Lambda_{1},\ldots,\Lambda_{q}$ be a partition of the index set $\{1,\ldots,s\}$, such that the cardinality of $\Lambda_{j}$ is $s_{j}$, and $s_{1}+\cdots+s_{q}=s$, so that the cycle $\pi_{j}$ acts on the $\Lambda_{j}$ and $f|_{\Lambda_{j}}:\bigvee_{\ell\in\Lambda_{j}}X_{\ell}\to\bigvee_{\ell\in\Lambda_{j}}X_{\ell}$, is a permutative map with permutation $\pi_{j}$. Hence $f=f_{\Lambda_{1}}\vee\cdots\vee f_{\Lambda_{q}}$ is the decomposition detailed in Lemma 5 below.
@@
We shall require an additional condition on the spaces $X_{i}$ for this class of maps. We say that a permutative self-map on $\bigvee_{i=1}^{s}X_{i}$ is squared by blocks if the spaces $X_{j}$, for $j\in\Lambda_{i}$, $1\leq i\leq q$, are pairwise homeomorphic. In the particular case that a permutative map associated with a cyclic permutation is squared by blocks, all the spaces $X_{j}$, for $1\leq j\leq s$ are homeomorphic.
In the following lines we introduce the Lefschetz numbers and the Lefschetz zeta function. Let $Y$ be a compact Euclidean neighborhood retract (ENR for short) of dimension $N$ (cf. [19, p. 527]). We recall that compact polyhedra and CW complexes are ENR-s, moreover the class of ENR is closed under wedge sums. Let $f:Y\to Y$ be a continuous map, and $f_{*k}:H_{k}(Y,\mathbb{Q})\to H_{k}(Y,\mathbb{Q})$ the induced maps on the $k$-rational homology group of $Y$, with $0\leq k\leq N$. The Lefschetz number $L(f)$ is defined as
@@
A classic result connecting algebraic topology with fixed point theory is the Lefschetz Fixed Point Theorem which establishes the existence of a fixed point if $L(f)\neq 0$ (cf. [24, 8]). As it is well-known the converse of this theorem is not true, consider for example the identity map on the torus.
Theorem 1 reduces the Lefschetz numbers and zeta function of the map $f$ in terms of the Lefschetz numbers and zeta functions of its coordinate maps.
@@
Let $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ be a continuous permutative squared by blocks self-map with associated permutation $\sigma$. Let $s=s_{1}+s_{2}+\cdots+s_{q}$ be the partition associated with the cyclic representation of $\sigma$. Then, for any partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of the set $\{1,2,\ldots,s\}$ for which $\Lambda_{j}=s_{j}$ for each $j=1,2,\ldots,q$, we have that
@@
[…]
@@
[…]
@@
We remark the fact $\ell(f^{m})\neq 0$ does not necessarily implies the existence of a periodic point of period $m$ for $f$. However, this holds in some particular situations. For instance, whenever the space $Y$ is a differentiable manifold and $f$ is a Morse-Smale diffeomorphism (see [30], for a survey on the subject) or a $C^{1}$ transversal map (cf. [25]). The terminology of the set of algebraic periods is mentioned explicitly in [14]. The intersection of the set of algebraic periods and the odd non-negative integers have been widely studied in the context of quasi–unipotent maps (in this case this intersection set is called the minimal set of Lefschetz periods) during the last years, see for example [30] and references therein.
@@
Let $f\colon\bigvee_{i=1}^{s}X_{i}\to\bigvee_{i=1}^{s}X_{i}$ be a continuous permutative squared by blocks self-map with associated permutation $\sigma$. Let $s=s_{1}+s_{2}+\cdots+s_{q}$ be the partition associated with the cyclic representation of $\sigma$. Then, for any partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of the set $\{1,2,\ldots,s\}$ for which $\Lambda_{j}=s_{j}$ for each $j=1,2,\ldots,q$, we have for $m>1$,
@@
A straightforward consequence of this theorem is the following corollary:
@@
An interesting application of Theorem 1 in the case of self-maps on wedge sums of tori is Theorem 4. We recall that a continuous map $f:X\to X$ is Lefschetz periodic point free if $L(f^{m})=0$, for all $m\geq 1$, i.e. $\zeta_{f}(t)\equiv 1$.
@@
Let $f:\bigvee_{i=1}^{s}\mathbb{T}^{n}\to\bigvee_{i=1}^{s}\mathbb{T}^{n}$ be as a continuous permutative map with non-zero eigenvalues then $f$ is not Lefschetz periodic point free.
@@ 3. Proofs of the main results and applications
Throughout the article we shall asume that the spaces $X_{j}$ are path-connected. Given a continuous map $f\colon\bigvee_{j=1}^{s}X_{j}\to\bigvee_{j=1}^{s}X_{j}$ we shall see in the next lemma that it is possible to find a decomposition of the sum $\bigvee_{j=1}^{s}X_{j}$ into subsums of the form $\bigvee_{j\in\Lambda}X_{j}$ which are invariant under the action of $f$, so that each summand is indecomposable in the sense that no nontrivial subsum is invariant under $f$. We recall that a set $Z\subset X$ is an invariant set for the map $g\colon X\to X$, if $g(Z)\subset Z$.
@@
Let us consider the graph $\mathcal{G}_{f}=(\mathcal{V}_{f},\mathcal{E}_{f})$ having as a set of vertices $\mathcal{V}_{f}=\{1,2,\ldots,s\}$ and edges defined by the condition $\{i,j\}\in\mathcal{E}_{f}$ if and only if there exist $x_{i}\in X_{i}$ and $x_{j}\in X_{j}$ with $x_{i},x_{j}\not=x_{0}$ (here $x_{0}$ is the base point) such that $f(x_{i})=x_{j}$ or $f(x_{j})=x_{i}$. An equivalent relation is defined by $i\sim_{f}j$ if and only if $i=j$ or there is a path in $\mathcal{G}_{f}$ joining $i$ and $j$. The equivalent classes form the required partition $\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{q}$ of $\{1,2,\ldots,s\}$. Indeed: the fact $i\in\Lambda_{k}$ and $f(X_{i})\cap X_{j}\not$$\emptyset$ implies $j\in\Lambda_{k}$, yields $f(\bigvee_{\ell\in\Lambda_{k}}X_{\ell})\subset\bigvee_{\ell\in\Lambda_{k}}X_{\ell}$. Similarly $f(\bigvee_{\ell\not\in\Lambda_{k}}X_{\ell})\subset\bigvee_{\ell\not\in\Lambda_{k}}X_{\ell}$, the indecomposibility is a consequence of the connectness of the classes of the graph $\mathcal{G}_{f}$. ∎
@@
Let us consider the example of the permutation map described before with $s=3$:
@@
where $O_{{ij}_{*k}}:H_{k}(X_{i},\mathbb{Q})\to H_{k}(X_{j},\mathbb{Q})$ is the null linear map, and $\sigma(1)=2$, $\sigma(2)=1$, $\sigma(3)=3$. In the case that $f$ is squared by blocks, we have $X_{1}$ is homeomorphic to $X_{2}$.
@@
In Theorem 1 we express the Lefschetz numbers and the Lefschetz zeta function of a squared by blocks permutative map on $\bigvee_{j=1}^{s}X_{j}$ in terms of its coordinate maps. The proof is divided in some Lemmas, that state the result first for diagonal maps, later for cyclic permutative maps, and finally we reduce the general case to these cases.
@@
We would like to point out that the statement (a) of Lemma 6 generalizes Theorem 3.2.1 of [11], in the context of wedge sums of spaces.
@@
[…]
@@
For the purpose of providing a formula that can be used for effective computations we can consider particular examples of the sets $\Lambda$’s. For example, if the orbits of $\sigma$ are segments of non-decreasing size, then they must begin at the points $1$, $_{1}, 1+s1+s$$_{1}+s_{2}$, $\ldots,1+s_{1}+\cdots+s_{q-1}$ and consequently have the form
@@
The Lefschetz zeta function of a continuous self-map $f:Y\to Y$, on any compact polyhedron $Y$, is a rational function. Moreover it satisfies the formula (cf. [12, pg. 38]):
@@
In Theorem 2, we express the Dold coefficients of a squared by blocks permutative map on $\bigvee_{j=1}^{s}X_{j}$ in terms of its coordinate maps. The strategy of the proof is similar to the one used in the proof of Theorem 1, i.e. we start with the case of diagonal maps (Lemma 10), later we consider cyclic permutative maps (Lemma 11) and finally these results are used for proving the general case.
@@
Part (a) is straightforward from Lemma 7-(a). For part (b), notice that if $r|m$ and $m\not\equiv 0\pmod{s}$ then $r\not\equiv 0\pmod{s}$. Lemma 7-(b), states $L(f^{r})=1$, for $r\not\equiv 0\pmod{s}$. Hence, for $m\not\equiv 0\pmod{s}$, it yields
@@
[…]
@@
4. Self-maps on wedge sums of tori
In this section we shall consider continuous map on a wedge sum of $n$-dimensional tori, i.e. $\bigvee_{i=1}^{s}\mathbb{T}^{n}$, with $n>1$. As we pointed out in the Introduction, when $n=1$, we have $\bigvee_{i=1}^{s}{\mathbb{S}}^{1}$ is a graph and graph maps have been extensively studied.